An electric utility is considering a new power plant in northern Arizona. Power from the plant would be sold in the Phoenix area, where it is badly needed. Because the firm has received a permit, the plant would be legal; but it would cause some air pollution. The company could spend an additional $40 million at Year 0 to mitigate the environmental problem, but it would not be required to do so. The plant without mitigation would cost $240.20 million, and the expected cash inflows would be $80 million per year for 5 years. If the firm does invest in mitigation, the annual inflows would be $84.78 million. Unemployment in the area where the plant would be built is high, and the plant would provide about 350 good jobs. The risk adjusted WACC is 18%. a.Calculate the NPV and IRR with mitigation b.Calculate the NPV and IRR without mitigation
Answer #1

NPV is the net present value of cash flows which can be calculated as, NPV = -C0+C1/(1+r)+C2/(1+r)2+.........+CT/(1+r)T

​​​​​​where, -C0= Cash outflow

C= Cash inflow

r = Discount rate

T= Time

Sol. NPV and IRR with mitigation

Cash outflow = $(240.20+40)million

= $ 280.20 million

Cash inflows = $84.78 million

NPV of the project with mitigation

= -280200000+ 84780000/(1+.18)1+84780000/(1+.18)2+84780000/(1+.18)3+84780000/(1+.18)4+84780000/(1+.18)5

=-$15,078,440.84

IRR is the discount rate that equates the cash outflows to the cash inflows and can be calculated as

0 = -C0+C1/(1+IRR)1+......CT/(1+IRR)T

IRR =15.601%

(b) NPV and IRR without mitigation

Cash outflows without mitigation = $240.20million

Cash inflows without mitigation = $80 million

NPV and IRR can be calculated with the formulas described above or with the financial calculator or excel

NPV without mitigation will be = $9,973,681.68

IRR will be = 19.82%